# S3Attention: Improving Long Sequence Attention with Smoothed Skeleton Sketching

### v1

@@ III-A Skeleton Sketching Based Attention
Thirdly, we resolve the error bound estimation. For the noisy matrix $\bm{X}+\bm{E}$, we directly apply [75, Corollary 4.3]. Thus, we have
@@ IV-C Transfer Learning in GLUE Tasks
We evaluate the transfer learning ability of the proposed model in the pretraining-finetuning paradigm in NLP tasks. We pre-train vanilla BERT [2], FNet [54], PoNet [55] and our S3Attention with the same MLM loss in [2] on English Wikitext-103 and BooksCorpus datasets. All models are uncased and pre-trained with the same configuration with 1 million steps at most. We report the best GLUE results for each model from multiple hyper-parameters configurations in Table 3, and the detailed training configurations in Table 15 in Table 3. Our S3Attention reaches 77.01 average scores (96.0% of the accuracy of vanilla BERT), which also outperform FNet by 4.6% and PoNet by 0.3% relatively.
@@ IV-F Model Parameters Impact
S3Attention introduces three extra hyperparameters, $r,s_{1}$ and $s_{2}$. We test the influence when varying them and report results in Table 6. We use S3Attention $(r,s_{1},s_{2}=8)$ as the baseline model and other parameters are reported in Table 10 in the supplementary material.
@@ IV-H Ablation Study
This subsection provides an ablation test on four components: Fourier Convolution, Convolution Stem, Column Attention, and Row Attention. We use S3Attention with ($r,s_{1},s_{2}=8$) as the baseline, and the settings are detailed in Table 11 in the supplementary material. In Table 7, we present the accuracy changes when removing each component. The performance-decreasing results indicate all four components of S3Attention are necessary to reach promising results. The most significant component is Column Attention which leads 8.28 average accuracy difference. It states that a good summary of the whole sequence is important. Similar observations are also reported in Transformer-LS [23] and XCiT [64], where the spirit of Attention over columns is used in the dynamic project and Cross-Covariance Attention respectively. The second most effective part is Fourier Convolution. It reaches a 13.89% accuracy difference in the Retrieval task involving two 4k sequences. Fourier Convolution also works well on shorter sequence tasks (e.g., Image and Pathfinder) and brings a 6.12% accuracy difference.
@@ V Concluding Remarks
In this section, we report the configurations in Tables 8-11 for the experiments in Section IV.
Table 8: Experiment Configuration of S3Attention $(r,s_{1},s_{2}=8)$.
@@
Table 9: Experiment Configuration of S3Attention (best).
@@
Table 10: Experiment Configuration for Model Parameters Impact.
@@
Table 11: Experiment Configuration for Ablation.
@@
We have already provided the average of five runs with different random seeds in Table 1. Here we also provide the standard deviations for these experiments in Table 12.
Table 12: Accuracy on Long Range Arena (LRA) with standard errors shown in parenthesis. All results are averages of 5 runs with different random seeds.
@@
Time series datasets:1) ETT [81] dataset contains two sub-datasets: ETT1 and ETT2, collected from two separated counties. Each of them has two versions of sampling resolutions (15min & 1h). ETT dataset contains multiple time series of electrical loads and one time sequence of oil temperature. 2) Electricity [^3] https://archive.ics.uci.edu/ml/datasets/ElectricityLoadDiagrams20112014. dataset contains the electricity consumption for more than three hundred clients with each column corresponding to one client. 3) Exchange [85] dataset contains the current exchange of eight countries. 4) Traffic [^4] http://pems.dot.ca.gov. dataset contains the occupation rate of freeway systems in California, USA. 5) Weather [^5] https://www.bgc-jena.mpg.de/wetter. dataset contains 21 meteorological indicators for a range of one year in Germany. 6) Illness [^6] https://gis.cdc.gov/grasp/fluview/fluportaldashboard.html. dataset contains the influenza-like illness patients in the United States. Table 13 summarizes all the features of the six benchmark datasets. They are all split into the training set, validation set and test set by the ratio of 7:1:2 during modeling.
Table 13: Details of time series benchmark datasets.
@@
In this section, we verify Fourier convolution component in the Smoother block can reduce the incoherence value in the early training stage. We use S3Attention with ($r,s_{1},s_{2}=8$) as the test model and test on an NLP dataset: Text, and a vision dataset: Pathfinder. We compute the $\mu$-incoherence value. of the token matrix before and after the Fourier convolution (denoted as $\mu_{\bm{X}}$ and $\mu_{\bm{X}^{\mathrm{smooth}}}$, respectively) for each sample in the validation dataset. Since we do not explicitly force the token matrix to be low-rank required by Definition 1, we report the incoherence value for different rankness settings ($\mathrm{rank}=16$ and $\mathrm{rank}=32$) approximately, and the mean and standard deviation of incoherence value can be found in Table 14. The average incoherence value was reduced 30% after the Fourier convolution in both datasets. Moreover, We observe that the standard deviation significantly decreases, which suggests the Fourier convolution may also potentially stabilize the training procedure.
Table 14: The average incoherence parameters after 100 training steps with standard errors shown in the parenthesis.
@@
Table 15: The training configurations for Pretraining and GLUE tasks

### v3

@@ III-A Skeleton Sketching Based Attention
Thirdly, we resolve the error bound estimation. For the noisy matrix $\bm{X}+\bm{E}$, we directly apply [74, Corollary 4.3]. Thus, we have
@@ IV-C Transfer Learning in GLUE Tasks
We evaluate the transfer learning ability of the proposed model in the pretraining-finetuning paradigm in NLP tasks. We pre-train vanilla BERT [2], FNet [54], PoNet [55] and our S3Attention with the same MLM loss in [2] on English Wikitext-103 and BooksCorpus datasets. All models are uncased and pre-trained with the same configuration with 1 million steps at most. We report the best GLUE results for each model from multiple hyper-parameters configurations in Table 3, and the detailed training configurations in Table S8 in Table 3. Our S3Attention reaches 77.01 average scores (96.0% of the accuracy of vanilla BERT), which also outperform FNet by 4.6% and PoNet by 0.3% relatively.
@@ IV-F Model Parameters Impact
S3Attention introduces three extra hyperparameters, $r,s_{1}$ and $s_{2}$. We test the influence when varying them and report results in Table 6. We use S3Attention $(r,s_{1},s_{2}=8)$ as the baseline model and other parameters are reported in Table S3 in the supplementary material.
@@ IV-H Ablation Study
This subsection provides an ablation test on four components: Fourier Convolution, Convolution Stem, Column Attention, and Row Attention. We use S3Attention with ($r,s_{1},s_{2}=8$) as the baseline, and the settings are detailed in Table S4 in the supplementary material. In Table 7, we present the accuracy changes when removing each component. The performance-decreasing results indicate all four components of S3Attention are necessary to reach promising results. The most significant component is Column Attention which leads 8.28 average accuracy difference. It states that a good summary of the whole sequence is important. Similar observations are also reported in Transformer-LS [23] and XCiT [64], where the spirit of Attention over columns is used in the dynamic project and Cross-Covariance Attention respectively. The second most effective part is Fourier Convolution. It reaches a 13.89% accuracy difference in the Retrieval task involving two 4k sequences. Fourier Convolution also works well on shorter sequence tasks (e.g., Image and Pathfinder) and brings a 6.12% accuracy difference.
@@ V Concluding Remarks
In this section, we report the configurations in Tables S1-S4 for the experiments in Section IV.
Table S1: Experiment Configuration of S3Attention $(r,s_{1},s_{2}=8)$.
@@
Table S2: Experiment Configuration of S3Attention (best).
@@
Table S3: Experiment Configuration for Model Parameters Impact.
@@
Table S4: Experiment Configuration for Ablation.
@@
We have already provided the average of five runs with different random seeds in Table 1. Here we also provide the standard deviations for these experiments in Table S5.
Table S5: Accuracy on Long Range Arena (LRA) with standard errors shown in parenthesis. All results are averages of 5 runs with different random seeds.
@@
Time series datasets:1) ETT [80] dataset contains two sub-datasets: ETT1 and ETT2, collected from two separated counties. Each of them has two versions of sampling resolutions (15min & 1h). ETT dataset contains multiple time series of electrical loads and one time sequence of oil temperature. 2) Electricity [^3] https://archive.ics.uci.edu/ml/datasets/ElectricityLoadDiagrams20112014. dataset contains the electricity consumption for more than three hundred clients with each column corresponding to one client. 3) Exchange [1] dataset contains the current exchange of eight countries. 4) Traffic [^4] http://pems.dot.ca.gov. dataset contains the occupation rate of freeway systems in California, USA. 5) Weather [^5] https://www.bgc-jena.mpg.de/wetter. dataset contains 21 meteorological indicators for a range of one year in Germany. 6) Illness [^6] https://gis.cdc.gov/grasp/fluview/fluportaldashboard.html. dataset contains the influenza-like illness patients in the United States. Table S6 summarizes all the features of the six benchmark datasets. They are all split into the training set, validation set and test set by the ratio of 7:1:2 during modeling.
Table S6: Details of time series benchmark datasets.
@@
In this section, we verify Fourier convolution component in the Smoother block can reduce the incoherence value in the early training stage. We use S3Attention with ($r,s_{1},s_{2}=8$) as the test model and test on an NLP dataset: Text, and a vision dataset: Pathfinder. We compute the $\mu$-incoherence value. of the token matrix before and after the Fourier convolution (denoted as $\mu_{\bm{X}}$ and $\mu_{\bm{X}^{\mathrm{smooth}}}$, respectively) for each sample in the validation dataset. Since we do not explicitly force the token matrix to be low-rank required by Definition 1, we report the incoherence value for different rankness settings ($\mathrm{rank}=16$ and $\mathrm{rank}=32$) approximately, and the mean and standard deviation of incoherence value can be found in Table S7. The average incoherence value was reduced 30% after the Fourier convolution in both datasets. Moreover, We observe that the standard deviation significantly decreases, which suggests the Fourier convolution may also potentially stabilize the training procedure.
Table S7: The average incoherence parameters after 100 training steps with standard errors shown in the parenthesis.
@@
Table S8: The training configurations for Pretraining and GLUE tasks
@@
Additional References
  • [1] G. Lai, W.-C. Chang, Y. Yang, and H. Liu, “Modeling long-and short-term temporal patterns with deep neural networks,” in The 41st International ACM SIGIR Conference on Research & Development in Information Retrieval, 2018, pp. 95–104.